# Thinking with Images for Multimodal Reasoning: Foundations, Methods, and Future Frontiers

### v2

@@ Abstract
Recent progress in multimodal reasoning has been significantly advanced by textual Chain-of-Thought (CoT), a paradigm where models conduct reasoning within language. This text-centric approach, however, treats vision as a static, initial context, creating a fundamental “semantic gap” between rich perceptual data and discrete symbolic thought. Human cognition often transcends language, utilizing vision as a dynamic mental sketchpad. A similar evolution is now unfolding in AI, marking a fundamental paradigm shift from models that merely think about images to those that can truly think with images. This emerging paradigm is characterized by models leveraging visual information as intermediate steps in their thought process, transforming vision from a passive input into a dynamic, manipulable cognitive workspace. In this survey, we chart this evolution of intelligence along a trajectory of increasing cognitive autonomy, which unfolds across three key stages: from external tool exploration, through programmatic manipulation, to intrinsic imagination. To structure this rapidly evolving field, our survey makes four key contributions. (1) We establish the foundational principles of the think with image paradigm and its three-stage framework. (2) We provide a comprehensive review of the core methods that characterize each stage of this roadmap. (3) We analyze the critical landscape of evaluation benchmarks and transformative applications. (4) We identify significant challenges and outline promising future directions. Through this structured overview, we aim to offer a clear roadmap for future research towards more powerful and human-aligned multimodal AI [^1] We maintain a real-time GitHub repository tracking progress at: https://github.com/zhaochen0110/Awesome_Think_With_Images..
@@ 1 Introduction
In this established paradigm, which we term “Thinking about Images”, the visual modality primarily serves as a static, initial context. LMM first “sees” an image, encodes it into a fixed set of features, and then operates exclusively within the textual domain to conduct its reasoning. The image is the premise, but language is the exclusive medium of thought. While powerful, this text-centric approach has exposed a fundamental limitation: the “semantic gap” between the rich, continuous, and often ambiguous nature of the visual world and the discrete, symbolic structure of language (Li et al., 2022). This initial, one-time encoding flattens the visual world into a static representation, creating a critical information bottleneck. Consequently, models often falter on tasks that demand deeper, iterative visual engagement, such as complex physical reasoning (Balazadeh et al., 2024), precise spatial manipulation (Gupta and Kembhavi, 2023), or long-horizon planning in interactive environments (Pahuja et al., 2025; Wang et al., 2025a).
@@
Now, we are in the midst of a profound sea change in multimodal reasoning. The community is moving beyond merely thinking about images with text, towards a new paradigm where models can truly think with images. This represents a fundamental shift in the cognitive role of vision: from a passive, fixed input to a dynamic, manipulable cognitive workspace. Much like a human using a sketchpad, models are now being empowered to actively query, modify, and even generate new visual information as integral, intermediate steps within their reasoning process. This ability to form a “visual chain of thought” is not a mere extension of textual CoT, but a revolutionary step towards a more holistic and human-like form of cognition (Larkin and Simon, 1987). We argue that this emerging paradigm, “Thinking with Images”, constitutes the next frontier for multimodal AI. It is defined by its core principle: to utilize visual representations as a form of manipulable and verifiable thought, thereby empowering models to actively see, manipulate, and reason with visual information as intermediate steps in a cognitive process. This survey provides the first comprehensive and systematic overview of this nascent and rapidly accelerating field.
@@
To structure this rapidly evolving landscape, we propose a conceptual framework that charts the progression of this paradigm through three stages of increasing cognitive autonomy. These stages represent the different mechanisms (the “How”) through which a model can achieve its visual reasoning goals (the “Why”): from acting as a “commander” that orchestrates external visual tools, to evolving into a “visual programmer” that generates code for bespoke operations, and ultimately becoming a “visual thinker” capable of intrinsic imagination and simulation. This three-stage evolution will be detailed in Section 2. The proliferation of methods across these stages naturally raises a critical question that this survey seeks to answer:
@@
This survey is organized to systematically address this question, following the structure of our proposed taxonomy illustrated in Figure 1. We begin by establishing the foundations of this paradigm in §2. We then delve into the methodologies of each three stages in §3, §4, and §5, respectively. Following this, we review the critical landscape of evaluation benchmarks and implementation frameworks in §6. We then explore the transformative applications of this paradigm in §7, and discuss prominent challenges and future directions in §8. By providing a clear taxonomy and a forward-looking perspective, we aim to not only synthesize existing knowledge but also to inspire future research towards building more powerful, intuitive, and truly multimodal AI.
@@
Manipulation (§4) [ Prompt-Based
@@
CoSyn (Yang et al., 2025a), etc. , leaf2, text width=32.7em ] ] [ RL-Based
@@
CoT-VLA (Zhao et al., 2025), T2I-R1 (Jiang et al., 2025a), BLIP3-o (Chen et al., 2025b), etc. , leaf, text width=32.7em ] ] [ RL-Based
@@
GeoMVerse (Kazemi et al., 2023), MATP-Bench (He et al., 2025b), VisioMath (Li et al., 2025b), MV-MATH (Wang et al., 2025f),
MathVista (Lu et al., 2024), ScienceQA (Lu et al., 2022), MMIE (Xia et al., 2024a), PuzzleBench (Zhang et al., 2025c), etc. , leaf5, text width=38.5em ] ] [ Implementation
@@
Unsloth (UnslothAI, 2023), etc. , leaf5, text width=38.5em ] ] ] [ Applications (§7) [ Interactive Systems
and User Interfaces
 (§7.1) [ e.g., CogAgent (Hong et al., 2024), MobileLMM (Liu et al., 2024b), ShowUI (Lin et al., 2024), UI-TARS (Qin et al., 2025),
Aguvis (Xu et al., 2024), Explorer (Pahuja et al., 2025), TongUI (Zhang et al., 2025d), PC Agent-E (He et al., 2025c), etc. , leaf4, text width=38.5em ] ] [ Embodied AI
and Robotics (§7.2) [ e.g., PaLM-E (Driess et al., 2023), Interleave-VLA (Fan et al., 2025b), VIMA (Jiang et al., 2023), CoT-VLA (Zhao et al., 2025),
Embodied ECoT (Zawalski et al., 2025), VPRL (Xu et al., 2025), Genie (Bruce et al., 2024), 3D-VLA (Zhen et al., 2024), etc. , leaf4, text width=38.5em ] ] [ AI for Science (§7.3) [ e.g., GeoCha (Kuckreja et al., 2023), PCBs (Balazadeh et al., 2024), MAPS (Zhu et al., 2025b), ChemMLLM (Tan et al., 2025), etc. , leaf4, text width=38.5em ] ] [ Healthcare and
Medicine (§7.4) [ e.g., HealthGPT (Lin et al., 2025), AOR (Li et al., 2025c), QoQ-Med (Dai et al., 2025a), MMedAgent-RL (Xia et al., 2025a), etc. , leaf4, text width=38.5em ] ] [ Education and
Training (§7.5) [ e.g., Interactive Sketchpad (Chen et al., 2025d), MathVista (Lu et al., 2023), DiagramGPT (Zala et al., 2024),
VidAAS (Lee et al., 2024), etc. , leaf4, text width=38.5em ] ] ] ]
Figure 1: The taxonomy of “Thinking with Images” organizes the field into three main categories: core methodologies (detailed across three stages of increasing intelligence), evaluation benchmarks and frameworks, and key applications. Representative works are listed at the leaves.
@@ 1.1 The Position of Our Survey
Earlier reviews have laid essential groundwork for understanding LMMs, yet they share a common perspective: vision is mainly treated as context while language remains the primary vehicle for reasoning. General-scope surveys such as Liang et al. (2024), Zhang et al. (2024c), and Wu et al. (2023a) catalogue architectures, pre-training corpora, and evaluation protocols, but only briefly touch on how a model might use visual information once it has been encoded. Xie et al. (2024a) extends the discussion to agent settings, yet still emphasises tool calling rather than internal visual cognition. Domain-specific reviews on mathematical reasoning (Yan et al., 2024), hallucination (Liu et al., 2024c), and benchmark (Li et al., 2025d) delve deeper into task details but inherit the same text-centric framing. Very recent work has begun to consider explicit reasoning mechanisms. Wang et al. (2025g) surveys multimodal chain-of-thought prompting, and Li et al. (2025e) discusses perception–reason–plan pipelines. Even here, the visual modality is largely passive: models describe or annotate an image once, then proceed with textual deliberation.
@@ 2.1 The Cognitive Imperative: Why We Need “Thinking with Images”?
The prevailing paradigm for multimodal reasoning, where a model first perceives a visual input and then performs subsequent reasoning exclusively through a textual CoT (Wei et al., 2022; Zhang et al., 2023a; Wang et al., 2025g), has encountered a fundamental ceiling. This approach is constrained by a “semantic gap” between the high-dimensional, continuous nature of the visual world and the discrete, abstract structure of language. When an image is encoded into a single, static feature vector, the rich, relational structure of the visual world is flattened, making the process of reasoning itself intractable. Consequently, even models with superhuman perception can appear cognitively brittle when facing problems that require more than a static glance. We illustrate this fundamental contrast in Figure 2 (left), where the “Thinking about Images” approach, limited to a single static perception, fails due to lossy, one-time encoding. The “Thinking with Images” paradigm, however, directly overcomes this limitation by transforming the image from a static input into a dynamic and manipulable cognitive workspace. This shift empowers models with a more robust form of cognition, which, as illustrated in Figure 2 (right), manifests through three key capabilities:
@@
  • Goal-Oriented Generative Planning. Finally, the third capability unlocks goal-oriented generative planning by transforming the model’s generative capacity into an engine for simulation. This is essential for tasks requiring physical intuition and forward planning. For example, to determine if a robot arm can grasp an apple without knocking over a nearby glass, a textual CoT must rely on fallible symbolic logic about coordinates and trajectories. In stark contrast, a model thinking with images can simulate the entire action by generating a hypothetical future frame. In one such imagined future, the model might depict the robot’s gripper colliding with the glass. This generated image is not merely an illustration; it is a form of perceptual self-critique. The visual inconsistency of the collision provides a direct, non-linguistic error signal, forcing the model to re-evaluate its plan. This process of visual simulation is powerful because it outsources validation to the coherence of the physical world itself. A physically implausible plan results in a visually nonsensical image, grounding the model’s reasoning in a way that abstract text never could.
@@ 2.2 How Does It Work? The Three-Stage Evolution
The realization of “Thinking with Images” is best understood as an evolutionary journey. We categorize the diverse implementation pathways into three principal stages. These stages reflect a general trend from relying on external systems to developing internalized cognitive functions. To illustrate this evolution, consider the complex task of planning how to rearrange a cluttered room to accommodate a new, larger sofa.
@@ 2.4 Unique Challenges
The dense nature of visual information causes errors to propagate in uniquely damaging ways. A flawed assertion in a textual chain can introduce a logical contradiction (Su et al., 2024; Jia et al., 2025; Sun et al., 2024b). However, A visual error corrupts the underlying perception of reality (Dosovitskiy et al., 2020). Consider a model using a zoom-in tool to identify a product’s material. If it erroneously focuses on the wooden table underneath, its subsequent analysis is not merely off-topic. Instead of simply failing, it may begin a new, internally coherent line of reasoning about a completely irrelevant subject, generating a plausible-sounding analysis of wood grain and varnish sheen. This initial mistake establishes a false ground truth for all subsequent deliberation, creating a foundational error in perception whose interconnected falsehoods poison the entire reasoning process.
@@
The “Thinking with Images” paradigm introduces a profound challenge: a single mode of visual thinking is inadequate for the diverse nature of visual tasks. Different problems demand different modes of visual thinking. For example, solving a geometry puzzle might require a constructive strategy of generating new visual elements like lines and circles (Wang et al., 2025d). Diagnosing a detail-rich image may need an analytical strategy involving a sequence of targeted zoom-ins and comparisons (Zheng et al., 2025). Navigating a maze could demand a simulative strategy where the model imagines future visual states (Li et al., 2025a). A model trained with one strategy will fail at tasks requiring another. Developing agents that can maintain a toolkit of diverse visual strategies and learn a meta-policy to select the appropriate one for a given task remains a nuanced and unsolved challenge.
Figure 3: The specific operational categories within each of the three “Thinking with Images” paradigms, illustrating the evolution from selection to creation and imagination. Stage 1: The model selects from a fixed toolkit, categorized into semantic grounding, perception exploration, and visual reasoning-aid tools. Stage 2: The model creates bespoke programs by composing visual, logical, and composite operations. Stage 3: The model generates internal thoughts, which can manifest as implicit latent reasoning, explicit visual reasoning, and interleaved multimodal reasoning.
@@ 3.2 Categories of Visual Tools
ZoomEye (Shen et al., 2024) LLaVA-v1.5-7/13B, etc.

### v3

@@ Abstract
Recent progress in multimodal reasoning has been significantly advanced by textual Chain-of-Thought (CoT), a paradigm where models conduct reasoning within language. This text-centric approach, however, treats vision as a static, initial context, creating a fundamental semantic gap between rich perceptual data and discrete symbolic thought. Human cognition often transcends language, utilizing vision as a dynamic mental sketchpad. A similar evolution is now unfolding in AI, marking a fundamental paradigm shift from models that merely think about images to those that can truly think with images. This emerging paradigm is characterized by models leveraging visual information as intermediate steps in their thought process, transforming vision from a passive input into a dynamic, manipulable cognitive workspace. In this survey, we chart this evolution of intelligence along a spectrum of increasing cognitive autonomy, which unfolds across three key stages: from external tool exploration, through programmatic manipulation, to intrinsic imagination. To structure this rapidly evolving field, our survey makes four key contributions. (1) We establish the foundational principles of the “Thinking with Images” paradigm and its three-stage framework. (2) We provide a comprehensive review of the core methods that characterize each stage of this roadmap. (3) We analyze the critical landscape of evaluation benchmarks and transformative applications. (4) We identify significant challenges and outline promising future directions. Through this structured overview, we aim to offer a clear roadmap for future research towards more powerful and human-aligned multimodal AI. [^1] We maintain a real-time GitHub repository tracking progress at: https://github.com/zhaochen0110/Awesome_Think_With_Images.
@@ 1 Introduction
In this established paradigm, which we term “Thinking about Images”, the visual modality primarily serves as a static, initial context. LMM first “sees” an image, encodes it into a fixed set of features, and then operates exclusively within the textual domain to conduct its reasoning. The image is the premise, but language is the exclusive medium of thought. While powerful, this text-centric approach has exposed a fundamental limitation: the semantic gap between the rich, continuous, and often ambiguous nature of the visual world and the discrete, symbolic structure of language (Li et al., 2022). This initial, one-time encoding flattens the visual world into a static representation, creating a critical information bottleneck. Consequently, models often falter on tasks that demand deeper, iterative visual engagement, such as complex physical reasoning (Balazadeh et al., 2024), precise spatial manipulation (Gupta and Kembhavi, 2023), or long-horizon planning in interactive environments (Pahuja et al., 2025; Wang et al., 2025a).
@@
Now, a new revolution is quietly unfolding in multimodal reasoning. Models are evolving beyond merely thinking about images with text, towards a new paradigm where they can truly think with images. This represents a fundamental shift in the cognitive role of vision: from a passive, fixed input to a dynamic, manipulable cognitive workspace. Much like a human using a sketchpad, models are now being empowered to actively query, modify, and even generate new visual information as integral, intermediate steps within their reasoning process. This ability to form a “visual chain of thought” is not a mere extension of textual CoT, but a revolutionary step towards a more holistic and human-like form of cognition (Larkin and Simon, 1987). We argue that this emerging paradigm, “Thinking with Images”, constitutes the next frontier for multimodal AI. It is defined by its core principle: to utilize visual representations as a form of manipulable and verifiable thought, thereby empowering models to actively see, manipulate, and reason with visual information as intermediate steps in a cognitive process. This survey provides the first comprehensive and systematic overview of this nascent and rapidly accelerating field.
@@
To structure this rapidly evolving landscape, we propose a conceptual framework that charts the progression of this paradigm through three stages of increasing cognitive autonomy. These stages represent the different mechanisms (the “How”) through which a model can achieve its visual reasoning goals (the “Why”): from acting as a “commander” that orchestrates external visual tools, to evolving into a “visual programmer” that generates code for tailored operations, and ultimately becoming a “visual thinker” capable of intrinsic imagination and simulation. This three-stage evolution will be detailed in Section 2. The proliferation of methods across these stages naturally raises a critical question that this survey seeks to answer:
@@
This survey is organized to systematically address this question, following the structure of our proposed taxonomy illustrated in Figure 1. We begin by establishing the foundations of this paradigm in §2. We then delve into the methodologies of each three stages in §3, §4, and §5, respectively. Following this, we review the critical landscape of evaluation benchmarks and implementation frameworks in §6. We then explore the transformative applications of this paradigm in §7, and discuss prominent challenges and future directions in §8. By providing a clear taxonomy and a forward-looking perspective, we aim to not only review existing knowledge but also to inspire future research towards building more powerful, intuitive, and truly multimodal AI.
@@
Manipulation (§4), font=, [ Prompt-Based
@@
CoSyn (Yang et al., 2025), etc. , leaf2, text width=32.7em ] ] [ RL-Based
@@
CoT-VLA (Zhao et al., 2025a), T2I-R1 (Jiang et al., 2025a), BLIP3-o (Chen et al., 2025b), etc. , leaf, text width=32.7em ] ] [ RL-Based
@@
GeoMVerse (Kazemi et al., 2023), VisioMath (Li et al., 2025b), MV-MATH (Wang et al., 2025f),
MathVista (Lu et al., 2024), ScienceQA (Lu et al., 2022), MMIE (Xia et al., 2024a), PuzzleBench (Zhang et al., 2025c), etc. , leaf5, text width=38.8em ] ] [ Implementation
@@
Unsloth (UnslothAI, 2023), etc. , leaf5, text width=38.8em ] ] ] [ Applications (§7) [ Interactive
Sys. and UI (§7.1), align=left [ e.g., CogAgent (Hong et al., 2024), MobileLMM (Liu et al., 2024b), ShowUI (Lin et al., 2024), UI-TARS (Qin et al., 2025),
Aguvis (Xu et al., 2024), Explorer (Pahuja et al., 2025), TongUI (Zhang et al., 2025d), PC Agent-E (He et al., 2025b), etc. , leaf4, text width=38.8em ] ] [ Embodied AI
and Robotics (§7.2), align=left [ e.g., PaLM-E (Driess et al., 2023), Interleave-VLA (Fan et al., 2025b), VIMA (Jiang et al., 2023), CoT-VLA (Zhao et al., 2025a),
Embodied ECoT (Zawalski et al., 2025), VPRL (Xu et al., 2025), Genie (Bruce et al., 2024), 3D-VLA (Zhen et al., 2024), etc. , leaf4, text width=38.8em ] ] [ AI for Science (§7.3), align=left, font=, [ e.g., GeoCha (Kuckreja et al., 2023), PCBs (Balazadeh et al., 2024), MAPS (Zhu et al., 2025b), ChemMLLM (Tan et al., 2025), etc. , leaf4, text width=38.8em ] ] [ Healthcare
and Medicine (§7.4), align=left [ e.g., HealthGPT (Lin et al., 2025), AOR (Li et al., 2025c), QoQ-Med (Dai et al., 2025a), MMedAgent-RL (Xia et al., 2025a), etc. , leaf4, text width=38.8em ] ] [ Education
and Training (§7.5), align=left [ e.g., Interactive Sketchpad (Chen et al., 2025d), DiagramGPT (Zala et al., 2024), VidAAS (Lee et al., 2024), etc. , leaf4, text width=38.8em ] ] ] ]
Figure 1: The taxonomy of the “Thinking with Images” paradigm, organized into three main branches: (1) core methodologies, evolving through Tool-Driven Visual Exploration, Programmatic Visual Manipulation, and Intrinsic Visual Imagination; (2) supporting benchmarks and frameworks; and (3) key applications. Representative works are listed on the leaves.
@@ 1.1 The Position of Our Survey
Earlier reviews have laid essential groundwork for understanding LMMs, yet they share a common perspective: vision is mainly treated as context while language remains the primary vehicle for reasoning. General-scope surveys such as Yin et al. (2024), Zhang et al. (2024c), and Wu et al. (2023a) catalogue architectures, pre-training corpora, and evaluation protocols, but only briefly touch on how a model might use visual information once it has been encoded. Xie et al. (2024a) extends the discussion to agent settings, yet still emphasises tool calling rather than internal visual cognition. Domain-specific reviews on mathematical reasoning (Yan et al., 2024), hallucination (Liu et al., 2024c), and benchmark (Li et al., 2025d) delve deeper into task details but inherit the same text-centric framing. Very recent work has begun to consider explicit reasoning mechanisms. Wang et al. (2025g) surveys multimodal CoT prompting, and Li et al. (2025e) discusses perception–reason–plan pipelines. Even here, the visual modality is largely passive: models describe or annotate an image once, then proceed with textual deliberation.
@@ 2.1 The Cognitive Imperative: Why We Need “Thinking with Images”?
The prevailing paradigm for multimodal reasoning, where a model first perceives a visual input and then performs subsequent reasoning exclusively through a textual CoT (Wei et al., 2022; Zhang et al., 2023a; Wang et al., 2025g), has encountered a fundamental ceiling. This approach is constrained by a semantic gap between the high-dimensional, continuous nature of the visual world and the discrete, abstract structure of language. When an image is encoded into a single, static feature vector, the rich, relational structure of the visual world is flattened, making the process of reasoning itself intractable. Consequently, even models with superhuman perception can appear cognitively brittle when facing problems that require more than a static glance. We illustrate this fundamental contrast in Figure 2 (left), where the “Thinking about Images” approach, limited to a single static perception, fails due to lossy, one-time encoding. The “Thinking with Images” paradigm, however, directly overcomes this limitation by transforming the image from a static input into a dynamic and manipulable cognitive workspace. This shift empowers models with a more robust form of cognition, which, as illustrated in Figure 2 (right), manifests through three key capabilities:
@@
  • Goal-Oriented Generative Planning. Finally, the third capability unlocks goal-oriented generative planning by transforming the model’s generative capacity into an engine for simulation. This is essential for tasks requiring physical intuition and forward planning. For example, to determine if a robot arm can grasp an apple without knocking over a nearby glass, a textual CoT must rely on fallible symbolic logic about coordinates and trajectories. In contrast, a model thinking with images can simulate the entire action by generating a hypothetical future frame. In one such imagined future, the model might depict the robot’s gripper colliding with the glass. This generated image is not merely an illustration; it is a form of perceptual self-critique. The visual inconsistency of the collision provides a direct, non-linguistic error signal, forcing the model to re-evaluate its plan. This process of visual simulation is powerful because it outsources validation to the coherence of the physical world itself. A physically implausible plan results in a visually nonsensical image, grounding the model’s reasoning in a way impossible by abstract text.
@@ 2.2 How Does It Work? The Three-Stage Evolution
The realization of “Thinking with Images” signifies an evolutionary journey. We categorize the diverse implementation pathways into three principal stages. These stages reflect a general trend from relying on external systems to developing internalized cognitive functions. To illustrate this evolution, consider the complex task of planning how to rearrange a cluttered room to accommodate a new, larger sofa.
@@ 2.4 Unique Challenges
The dense nature of visual information causes errors to propagate in uniquely damaging ways. A flawed assertion in a textual chain can introduce a logical contradiction (Su et al., 2024; Jia et al., 2025; Sun et al., 2024b). However, a visual error corrupts the underlying perception of reality (Dosovitskiy et al., 2020). Consider a model using a zoom-in tool to identify a product’s material. If it erroneously focuses on the wooden table underneath, its subsequent analysis, instead of simply failing, may begin a new, internally coherent line of reasoning about a completely irrelevant subject, generating a plausible-sounding analysis of wood grain and varnish sheen. This initial mistake establishes a false ground truth for all subsequent deliberation, creating a foundational error in perception whose interconnected falsehoods poison the entire reasoning process.
@@
The “Thinking with Images” paradigm introduces a profound challenge: a single mode of visual thinking is inadequate for the diverse nature of visual tasks. Different problems demand different modes of visual thinking. For example, solving a geometry puzzle might require a constructive strategy of generating new visual elements like lines and circles (Wang et al., 2025d). Diagnosing a detail-rich image may need an analytical strategy involving a sequence of targeted zoom-ins and comparisons (Zheng et al., 2025). Navigating a maze could demand a simulative strategy where the model imagines future visual states (Li et al., 2025a). A model trained with one strategy will fail at tasks requiring another. Developing agents that can maintain a toolkit of diverse visual strategies and learn a meta-policy to select the appropriate one for a given task remains a complex and unsolved challenge.
Figure 3: The specific operational categories within each of the three “Thinking with Images” paradigms, illustrating the evolution from selection to creation and imagination. Stage 1: The model selects from a fixed toolkit, categorized into semantic grounding, perception exploration, and visual reasoning-aid tools. Stage 2: The model creates customized programs by composing visual, logical, and composite operations. Stage 3: The model generates internal thoughts, which can manifest as implicit latent reasoning, explicit visual reasoning, and interleaved multimodal reasoning.
@@ 3.2 Categories of Visual Tools
ZoomEye (Shen et al., 2024) LLaVA-v1.5-7B, etc.
Chain-of-Spot (Liu et al., 2024d) LLaVA-v1.5-7B, etc.